### v2

# A note on the triple product property for finite groups with abelian normal subgroups of prime index

Abstract.
Three non-empty subsets $S,T,U$ of a group $G$ are said to satisfy the triple product property (TPP) if, for elements $s,s^{\prime}\in S$, $t,t^{\prime}\in T$, $u,u^{\prime}\in U$, the equation $s^{\prime}s^{-1}t^{\prime}t^{-1}u^{\prime}u^{-1}=1$ holds if and only if $s=s^{\prime}$, $t=t^{\prime}$, $u=u^{\prime}$. In this case $(S,T,U)$ is called a TPP triple of $G$ and $|S||T||U|$ is called the size of the triple. If $G$ is a finite group the triple product ratio of $G$ can be defined as the quantity $\rho(G):=\frac{\beta(G)}{|G|}$, where $\beta(G)$ is the largest size of a TPP triple of $G$, and a special case of this, the subgroup triple product ratio, is the quantity $\rho_{0}(G):=\frac{\beta_{0}(G)}{|G|}$, where $\beta_{0}(G)$ is the largest size of a TPP triple of $G$ composed only of subgroups. There is a conjecture that $\rho(G)\leq\frac{4}{3}$ if $G$ contains a cyclic subgroup of index $2$ [4, Conjecture 7.6]. This note proves a version of this conjecture for subgroups by showing that $\rho_{0}(G)\leq\frac{p^{2}}{2p-1}$ if $G$ is any finite group which contains an abelian normal subgroup of prime index $p$.
Key words and phrases:
fast matrix multiplication, triple product property, finite groups, TPP triple, TPP capacity, subgroup TPP capacity, TPP ratio, subgroup TPP ratio
2020 Mathematics Subject Classification:
Primary 20D60, Secondary 68R05
The author is not affiliated with any institution. The work contained in this note was originally completed in 2013-2014, and for this the author is grateful for the assistance of Prof. Sarah Hart, Birkbeck, University of London, and Dr. Peter M. Neumann, The Queen’s College, Oxford.
1. Introduction
Notation: Standard set-theoretic and group-theoretic notation is used. Groups will generally be finite, unless otherwise stated.
Definition 1.1.
[3, Definition 2.1] Let $G$ be a group, finite or infinite, and $S,T,U$ non-empty subsets of $G$ with cardinalities $|S|,|T|,|U|$ respectively. The triple $(S,T,U)$ is said to satisfy the triple product property (TPP) if
(1.1) $s^{\prime}s^{-1}t^{\prime}t^{-1}u^{\prime}u^{-1}=1\Longrightarrow s=s^{\prime},t=t^{\prime},u=u^{\prime},$
for all $s,s^{\prime}\in S$, $t,t^{\prime}\in T$, $u,u^{\prime}\in U$. In this case, $G$ is said to realise a TPP triple of parameter type, or simply type, $(|S|,|T|,|U|)$, and $|S|,|T|,|U|$ are called the parameters of the triple and the product $|S||T||U|$ the size of the triple. If, additionally, $S,T,U$ are subgroups of $G$ then $(S,T,U)$ is called a subgroup TPP triple of $G$, in which case the defining relation above simplifies to
(1.2) $stu=1\Longrightarrow s=t=u=1,$
for all $s\in S$, $t\in T$, $u\in U$.
TPP triples were first introduced by H. Cohn and C. Umans in 2003 to study the complexity of fast matrix multiplication [3]. Further applications and developments of this approach were described in [2, 4, 5]. However this note pursues a combinatorial problem for TPP triples in finite groups introduced by Peter M. Neumann in [6]. This involves two quantities. The first quantity, called the subgroup TPP ratio of $G$, is defined as
(1.3) $\rho_{0}(G):=\frac{\beta_{0}(G)}{|G|},$
where $\beta_{0}(G)$ is the subgroup TPP capacity of $G$ defined as
(1.4) $\beta_{0}(G):=\max\;\{|S||T||U|\mid(S,T,U)\textrm{ is a subgroup TPP triple of }G\}.$
The second quantity, called the TPP ratio of $G$, is defined as
(1.5) $\rho(G):=\frac{\beta(G)}{|G|},$
where $\beta(G)$ is the TPP capacity of $G$ defined as
(1.6) $\beta(G):=\max\;\{|S||T||U|\mid(S,T,U)\textrm{ is a TPP triple of }G\}.$
As subgroup TPP triples are special cases of TPP triples which are composed only of subgroups, clearly $\beta_{0}(G)\leq\beta(G)$ and $\rho_{0}(G)\leq\rho(G)$. Note that $G$ always realises the trivial subgroup TPP triple $(G,\{1\},\{1\})$, so that $\beta_{0}(G)\geq|G|$, or equivalently, $\rho_{0}(G)\geq 1$.
It is natural to seek best upper bounds for $\beta(G)$ and $\beta_{0}(G)$ (equivalently, for $\rho(G)$ and $\rho_{0}(G)$). It was shown by Cohn and Umans that if $G$ is a dihedral group then $\rho(G)\geq\frac{4}{3}$ [3]. P. Neumann derived a general upper bound for $\beta(G)$ [6, Corollary 3.2] that
(1.7) $\beta(G)\leq\left(\frac{1+\sqrt{1+8|G|}}{4}\right)^{3}<|G|^{3/2}$
using the following observation [6, Observation 3.1]:
Observation 1.2.
If $(S,T,U)$ is a TPP triple of a group $G$ then $|S|(|T|+|U|-1)\leq|G|$.
In 2012 Hedtke and Murthy conjectured [4, Table 1 and Conjectures 7.5-7.6], based on tables of TPP triple data in groups with cyclic subgroups of index $2$ of order up to $32$, obtained via exhaustive search algorithms in these groups implemented using the GAP computer algebra system, that
Conjecture 1.3.
If $G$ is a group with a cyclic subgroup of index $2$ then $\rho(G)\leq\frac{4}{3}$.
The author is aware of a proof (from a private communication from another researcher) of this conjecture for dihedral groups $D_{2n}$, i.e. that $\rho(D_{2n})\leq\frac{4}{3}$.
This note focuses on subgroup TPP ratio $\rho_{0}$ (equivalently, subgroup TPP capacity $\beta_{0}$), and proves a generalisation of the conjecture for $\rho_{0}$ for all groups with abelian normal subgroups of prime index. First, a few relevant properties and characterisations of TPP triples are stated, followed by proofs of two technical results needed to establish the main result.
2. Elementary Properties of TPP Triples
Let $G$ be a group.
Definition 2.1.
[6, p. 234] A TPP triple $(S,T,U)$ of $G$ is called basic if
(2.1) $S\cap T\cap U=\{1\}.$
For non-empty subsets $X,Y\subseteq G$ the quotient set $Q(X,Y)$ of $X$ and $Y$ is the set $Q(X,Y)=XY^{-1}=\{xy^{-1}|x\in X,y\in Y\}$. If $X=Y$ then $Q(X,Y)=Q(X)=XX^{-1}$.
Theorem 2.2.
[4, Theorem 3.1] For non-empty subsets $S,T,U\subseteq G$ the triple $(S,T,U)$ is a TPP triple if and only if
(2.2) $Q(S)\cap Q(T)Q(U)=Q(T)\cap Q(U)=\{1\},$
And if $S,T,U$ happen to be subgroups then the defining relation above simplifies to
(2.3) $S\cap TU=T\cap U=\{1\}.$
The following result states two invariance properties of TPP triples under permutations or certain kinds of set translations of its members.
Observation 2.3.
Let $(S,T,U)$ be any TPP triple of $G$, not necessarily basic in the sense defined above.
  1. (1)
    If $\pi\in\textrm{Sym}\{S,T,U\}$, i.e. $\pi$ is a permutation of $\{S,T,U\}$, then $(S^{\pi},T^{\pi},U^{\pi})$ is a TPP triple of $G$ (permutation invariance) [3, Lemma 2.1].
  2. (2)
    If $a,b,c,d\in G$ are any elements then $(dSa,dTb,dUc)$ is a TPP triple of $G$ (translation invariance) [6, Observation 2.1].
The translation invariance property for TPP triples means that any non-basic TPP triple can be translated to a basic TPP triple of the same type and size as the original triple. To be precise, if $(S,T,U)$ is any non-basic TPP triple of $G$, that is, when $1\notin S\cap T\cap U$, then elements $s\in S$, $t\in T$, $u\in U$ can be chosen such that the right-translated triple $(Ss^{-1},Tt^{-1},Uu^{-1})$ is a basic TPP triple of $G$ with parameters $|Ss^{-1}|=|S|$, $|Tt^{-1}|=|T|$, $|Uu^{-1}|=|U|$. This means that generally only basic TPP triples need be considered. Of course, subgroup TPP triples are necessarily basic.
All future references to TPP triples will be to basic TPP triples, unless otherwise stated.
The following two results are useful to note.
Observation 2.4.
Given a TPP triple $(S,T,U)$ of a group $G$ and a subgroup $H\leq G$ there is a subgroup TPP triple $(S\cap H,T\cap H,U\cap H)$ of $H$.
Proposition 2.5.
[3, Lemma 3.1] Let $(S,T,U)$ be a TPP triple of $G$.
  1. (1)
[…]
  2. (2)
    If $G$ is abelian then the mapping $(s,t,u)\longmapsto s^{-1}tu$ on $S\times T\times U$ is injective into $G$, and $|S||T||U|\leq|G|$.
3. Technical Results
Some technical results needed for the main result are stated and proved. These are based on the idea of decomposing a TPP triple of a given group into smaller TPP triples obtained by independently decomposing the members of the triple with the left (or right) cosets of a suitable subgroup, using the set-closure property of TPP triples in relation to subgroups (see Observation 2.4).
Definition 3.1.
Let $G$ be a group, $H\leq G$ and $G/H=\{gH\mid g\in G\}$ the quotient group of $G$ by $H$ whose order is denoted by $|G:H|$ (also, $[G:H]$). For a non-empty subset $S\subseteq G$ let the $H$-support of $S$ be the set $\overline{S}:=\{gH\in G/H\mid S\cap gH\neq\emptyset\}\subseteq G/H$, i.e. the set of all cosets $gH\in G/H$ which intersect with $S$. If $(S,T,U)$ is a TPP triple of $G$ let its $H$-restriction be the triple $(S_{0},T_{0},U_{0})$, where $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, $U_{0}:=U\cap H$, and this is a TPP triple of $H$.
Observation 3.2.
Let $G$ be a group and $H\leq G$.
(1) If $S\leq G$ and $\overline{S}$ is its $H$-support, then $|\overline{S}|=|S:S\cap H|$.
(2) If $S\leq G$ and $H\trianglelefteq G$, i.e. $H$ is normal in $G$, then $\overline{S}\leq G/H$.
(3) If $H$ is abelian and $(S,T,U)$ is a subgroup TPP triple of $G$ then $|S||T||U|\leq\frac{\sigma\tau\upsilon}{n}|G|$, where $|G:H|=n$, and $\sigma=|S:S\cap H|$, $\tau=|T:T\cap H|$, $\upsilon=|U:U\cap H|$.
Proof.
Let $G$ and $H$ be given as above and $|G:H|=n$.
(1) Let $S\leq G$ and $\overline{S}$ be its $H$-support as defined above. Let $gH\in\overline{S}$ and $S_{g}:=S\cap gH$. By definition $S_{g}\neq\emptyset$. If $x\in S_{g}$ then $|S_{g}|=|S\cap gH|=|xS\cap xH|=|x(S\cap H)|=|S\cap H|$, where $x(S\cap H)$ is a coset of $S\cap H\leq S$. As the cosets $gH\in G/H$ are disjoint so are the sets $S_{g}$, which thus partition $S$. If $\sigma=|\overline{S}|$ then $|S|=\sigma|S\cap H|$, i.e. $\sigma=|S:S\cap H|$. The same result holds if left cosets are replaced by right cosets.
(2) Let $S\leq G$. If $H\trianglelefteq G$ then using the Second Isomorphism Theorem $SH\leq G$ and $\overline{S}=SH/H\leq G/H$, where $SH/H\cong S/(S\cap H)$
(3) Let $(S,T,U)$ be a subgroup TPP triple of $G$. Then by Observation 2.4 $(S\cap H,T\cap H,U\cap H)$ is a subgroup TPP triple of $H$. If $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, $U_{0}:=U\cap H$ then by part (1) above $\sigma=|S:S_{0}|$, $\tau=|T:T_{0}|$, $\upsilon=|U:U_{0}|$, and, as by assumption $H$ is abelian, by Proposition 2.5 it follows that $|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{\sigma\tau\upsilon}\leq|H|$, which shows that $|S||T||U|\leq\sigma\tau\upsilon|H|=\frac{\sigma\tau\upsilon}{n}|G|$. ∎
Lemma 3.3.
Let $G$ be a group, $H\trianglelefteq G$ an abelian subgroup, $G/H=\{gH\mid g\in G\}$ the quotient group of $G$ by $H$, and $(S,T,U)$ a subgroup TPP triple of $G$. Let $\overline{S},\overline{T},\overline{U}\leq G/H$ be the $H$-supports of $S,T,U$ respectively, and $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, $U_{0}:=U\cap H$.
(1) If $xH\in\overline{S}\cap\overline{T}$ and $yH\in\overline{S}\cap\overline{U}$ then $S_{x}^{-1}T_{x}U_{0}$ and $S_{y}^{-1}U_{y}T_{0}$ are cosets of $S_{0}T_{0}U_{0}$ in $H$, where $S_{x}:=S\cap xH$, $T_{x}:=T\cap xH$, $U_{y}:=U\cap yH$. Additionally these cosets are non-trivial if $xH,yH\neq H$ and distinct if $xH\neq yH$.
(2) If $U_{0}=U\cap H\trianglelefteq G$ and $xH,yH\in\overline{S}\cap\overline{T}$ such that $xH\neq yH$ then $S_{x}^{-1}T_{x}U_{0}$, $S_{y}^{-1}T_{y}U_{0}$ are distinct cosets of $S_{0}T_{0}U_{0}$ in $H$, where $S_{x}:=S\cap xH$, $T_{x}:=T\cap xH$.
Proof.
Let $H$ be as given. We may assume $H<G$, otherwise there is nothing to prove. Note that for any $gH=Hg\in G/H$ its inverse is $(gH)^{-1}=g^{-1}H=Hg^{-1}$, and if $X_{g}\subseteq gH$ is any non-empty subset its inverse is $X_{g}^{-1}\subseteq g^{-1}H=Hg^{-1}$. Recall that $(S_{0},T_{0},U_{0})$ is a subgroup TPP triple of $H$, and, as $H$ is abelian, $|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|\leq|H|$. Also recall that, as subgroups with the TPP, $S,T,U$ satisfy $S\cap TU=T\cap U=\{1\}$.
(1) Let $S_{x},T_{x}$ be defined as above. As $S_{x},T_{x}\neq\emptyset$ elements $s_{x}\in S_{x}$, $t_{x}\in T_{x}$ can be chosen. Then $s_{x}^{-1}t_{x}\in H$ and $s_{x}^{-1}t_{x}(S_{0}T_{0}U_{0})=S_{0}s_{x}^{-1}t_{x}T_{0}U_{0}=S_{x}^{-1}T_{x}U_{0}\subseteq H$. This shows that $S_{x}^{-1}T_{x}U_{0}$ is a coset of $S_{0}T_{0}U_{0}$ in $H$, and $|S_{x}^{-1}T_{x}U_{0}|=|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|\leq|H|$. Clearly if $xH\neq H$, that is, if $x\in G\backslash H$, then $S_{x}^{-1}T_{x}U_{0}$ is a non-trivial coset of $S_{0}T_{0}U_{0}$ in $H$. Now, also assume that $yH\in\overline{S}\cap\overline{U}$ and let $S_{y}:=S\cap yH$ and $U_{y}:=U\cap yH$. By transposing $T$ and $U$, and $x$ and $y$, in the argument above, it can be shown that $S_{y}^{-1}U_{y}T_{0}$ is a coset of $S_{0}T_{0}U_{0}$ in $H$, and, furthermore, a non-trivial coset if $yH\neq H$.
Assume $xH\neq yH$. Suppose $S_{x}^{-1}T_{x}U_{0}\cap S_{y}^{-1}U_{y}T_{0}\neq\emptyset$. Then there are elements $s_{x}\in S_{x}$, $s_{y}\in S_{y}$, $t_{x}\in T_{x}$, $u_{y}\in U_{y}$, $t_{0}\in T_{0}$, and $u_{0}\in U_{0}$ such that $s_{x}^{-1}t_{x}u_{0}=s_{y}^{-1}u_{y}t_{0}$. We can rewrite this as $s_{x}s_{y}^{-1}u_{y}t_{0}u_{0}^{-1}t_{x}^{-1}=1$. Since $t_{0},u_{0}$ commute (as elements of the abelian group $H$), this can be rewritten as $s_{x}s_{y}^{-1}u_{y}u_{0}^{-1}t_{0}t_{x}^{-1}=1$. The TPP implies that $u_{y}=u_{0}$ and $t_{x}=t_{0}$, a contradiction. So $S_{x}^{-1}T_{x}U_{0}\cap S_{y}^{-1}U_{y}T_{0}=\emptyset$, which proves part (1).
(2) Assume $U_{0}=U\cap H\trianglelefteq G$ and $xH,yH\in\overline{S}\cap\overline{T}$ such that $xH\neq yH$. From above $S_{x}^{-1}T_{x}U_{0},S_{y}^{-1}T_{y}U_{0}\subseteq H$ are distinct cosets of $S_{0}T_{0}U_{0}$ in $H$. Suppose there are elements $s_{x}\in S_{x}$, $s_{y}\in S_{y}$, $t_{x}\in T_{x}$, $t_{y}\in T_{y}$, $u_{0},\widetilde{u}_{0}\in U_{0}$ such that $s_{x}^{-1}t_{x}u_{0}=s_{y}^{-1}t_{y}\widetilde{u}_{0}$. We can rewrite this as $st_{y}\bar{u}_{0}t_{x}^{-1}=1$, where $s=s_{x}s_{y}^{-1}$ and $\bar{u}_{0}=\widetilde{u}_{0}u_{0}^{-1}$. By the normality of $U_{0}$ in $G$ there is a $\hat{u}_{0}\in U_{0}$ such that $\bar{u}_{0}t_{x}^{-1}=t_{x}^{-1}\hat{u}_{0}$. So $st_{y}\bar{u}_{0}t_{x}^{-1}=st_{y}t_{x}^{-1}\hat{u}_{0}=1$. The TPP implies that $t_{x}=t_{y}$, a contradiction. Thus $S_{x}^{-1}T_{x}U_{0}$, $S_{y}^{-1}T_{y}U_{0}$ are distinct cosets of $S_{0}T_{0}U_{0}$ in $H$, which proves part (2). ∎
Theorem 3.4.
If $G$ is a group with an abelian normal subgroup of prime index $p$ then $\rho_{0}(G)\leq\frac{p^{2}}{2p-1}$.
Proof.
Let $G$ and $p$ be as given. Let $H\unlhd G$ be abelian of index $p$. Then $G/H$ is a cyclic group of order $p=|G:H|=\frac{|G|}{|H|}$ whose elements are $p$ (left) cosets $H_{0}=H,H_{1},\ldots,H_{p-1}$ of $H$ in $G$. Let $(S,T,U)$ be a subgroup TPP triple of $G$ with $H$-supports $\overline{S},\overline{T},\overline{U}$. By Observation 3.1 part (2) $\overline{S},\overline{T},\overline{U}\leq G/H$, and if $\sigma=|\overline{S}|$, $\tau=|\overline{T}|$, $\upsilon=|\overline{U}|$ then $\sigma,\tau,\upsilon\in\{1,p\}$. Let $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, $U_{0}:=U\cap H$. Then $(S_{0},T_{0},U_{0})$ is a subgroup TPP triple of $H$ (by Observation 2.4), and $S_{0}T_{0}U_{0}\mathrel{\unlhd}H$ (as $S_{0}T_{0}U_{0}\leq H$ and $H$ is abelian) and $|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{\sigma\tau\upsilon}\leq|H|$ and $|S||T||U|\leq\frac{\sigma\tau\upsilon}{p}|G|$ (by Proposition 2.5 and Observation 3.2).
Assume $|S||T||U|=\sigma\tau\upsilon|S_{0}||T_{0}||U_{0}|>|G|=p|H|$. Clearly at least two of $\sigma,\tau,\upsilon$ exceed $1$, that is, are equal to $p$. Without loss of generality we may assume $\sigma=\tau=p$. If $\upsilon=1$ then $|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{p^{2}}>\frac{|H|}{p}$, and this implies that $|S_{0}||T_{0}||U_{0}|=|H|$ , that is, $S_{0}T_{0}U_{0}=H$. However, $S$ and $T$ are supported on all $p$ cosets of $H$ in $G$, and using Lemma 3.3 any non-trivial coset $xH$ can be chosen such that $S_{x}:=S\cap xH\neq\emptyset$, $T_{x}:=T\cap xH\neq\emptyset$ and $S_{x}^{-1}T_{x}U_{0}$ is a non-trivial coset of $S_{0}T_{0}U_{0}$ in $H$, which is a contradiction. Thus $\sigma=\tau=\upsilon=p$ and $S,T,U\not\leq H$.
So the assumption $|S||T||U|>|G|=p|H|$ implies that $S_{0}T_{0}U_{0}$ is a subgroup of $H$ of order $|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{p^{3}}\leq|H|$. Furthermore applying part (1) of Lemma 3.3 shows that $\{S_{x}^{-1}T_{x}U_{0}\}_{xH\in G/H}$ and $\{S_{y}^{-1}U_{y}T_{0}\}_{yH\in G/H}$ are collections of cosets of $S_{0}T_{0}U_{0}$ in $H$ having only the trivial coset $S_{0}T_{0}U_{0}$ in common. As $H$ is maximal in $G$ and $U\not\leq H$ it follows that $U_{0}=U\cap H\trianglelefteq\langle U,H\rangle=G$. Now applying part (2) of Lemma 3.3 it follows that $\{S_{x}^{-1}T_{x}U_{0}\}$ is a collection of $p$ cosets of $S_{0}T_{0}U_{0}$ in $H$, and similarly that $T_{0}\trianglelefteq G$ and $\{S_{y}^{-1}U_{y}T_{0}\}_{yH\in G/H}$ is a collection of $p$ cosets of $S_{0}T_{0}U_{0}$ in $H$, these collections only intersecting trivially in $S_{0}T_{0}U_{0}$. This means that $S_{0}T_{0}U_{0}$ has at least (2p1)(2p-1) cosets in $H$, i.e. $|H:S_{0}T_{0}U_{0}|=\frac{|H|}{|S_{0}||T_{0}||U_{0}|}\geq 2p-1$, which means that $|S||T||U|=p^{3}|S_{0}||T_{0}||U_{0}|\leq p^{3}\cdot\frac{|H|}{(2p-1)}=p^{3}\cdot\frac{1}{p(2p-1)}|G|$$\frac{p^{2}}{2p-1}|G|$.