### v5

# A note on the triple product property for finite groups with abelian normal subgroups of prime index

Abstract.
Three non-empty subsets $S,T,U$ of a group $G$ are said to satisfy the triple product property (TPP) if, for elements $s,s^{\prime}\in S$, and $t,t^{\prime}\in T$, and $u,u^{\prime}\in U$, the equation $s^{\prime}s^{-1}t^{\prime}t^{-1}u^{\prime}u^{-1}=1$ holds if and only if $s=s^{\prime}$, $t=t^{\prime}$, $u=u^{\prime}$. If this is the case then $(S,T,U)$ is called a TPP triple of $G$ and $|S||T||U|$ the size of the triple. If $G$ is a finite group the triple product ratio of $G$ can be defined as the quantity $\rho(G):=\frac{\beta(G)}{|G|}$, where $\beta(G)$ is the largest size of a TPP triple of $G$, and a special case of this, the subgroup triple product ratio, is the quantity $\rho_{0}(G):=\frac{\beta_{0}(G)}{|G|}$, where $\beta_{0}(G)$ is the largest size of a TPP triple of $G$ composed only of subgroups. There is a conjecture that $\rho(G)\leq\frac{4}{3}$ if $G$ contains a cyclic subgroup of index $2$ [5, Conjecture 7.6]. This note proves a more general version of this conjecture for subgroups by showing that $\rho_{0}(G)\leq\frac{p^{2}}{2p-1}$ if $G$ is any finite group which contains an abelian normal subgroup of prime index $p$, and discusses its implications for $\rho$ for groups with cyclic normal subgroups of prime index, based on the known data for $\rho$ in such groups of small order.
Key words and phrases:
triple product property, finite groups, TPP triple, TPP capacity, subgroup TPP capacity, TPP ratio, subgroup TPP ratio, fast matrix multiplication
2020 Mathematics Subject Classification:
Primary 20D60, Secondary 68R05
The author is unaffiliated. With the exception of some minor updates, all the work contained in this note was originally completed in 2013-2014, and for this the author is grateful for the assistance of Prof. Sarah Hart, Birkbeck, University of London, and Dr. Peter M. Neumann, The Queen’s College, Oxford.
1. Introduction
Notation: Standard set-theoretic and group-theoretic notation is used. Groups will generally be finite, unless otherwise stated.
Definition 1.1.
[3, Definition 2.1] Let $G$ be a group, finite or infinite, and $S,T,U$ non-empty subsets of $G$ with cardinalities $|S|,|T|,|U|$ respectively. The triple $(S,T,U)$ is said to satisfy the triple product property (TPP) if
(1.1) $s^{\prime}s^{-1}t^{\prime}t^{-1}u^{\prime}u^{-1}=1\Longrightarrow s=s^{\prime},t=t^{\prime},u=u^{\prime},$
for all $s,s^{\prime}\in S$, $t,t^{\prime}\in T$, $u,u^{\prime}\in U$. In this case, $G$ is said to realise a TPP triple of parameter type, or simply type, $(|S|,|T|,|U|)$, and $|S|,|T|,|U|$ are called the parameters of the triple and the product $|S||T||U|$ the size of the triple. If, additionally, $S,T,U$ are subgroups of $G$ then $(S,T,U)$ is called a subgroup TPP triple of $G$, in which case the defining relation above simplifies to
(1.2) $stu=1\Longrightarrow s=t=u=1,$
for all $s\in S$, $t\in T$, $u\in U$.
TPP triples were first introduced by H. Cohn and C. Umans in 2003 to study the complexity of fast matrix multiplication [3]. Further applications and developments of this approach were described in [2, 5, 6]. However this note pursues the combinatorial problems for TPP triples in finite groups introduced by Peter M. Neumann in [7]. These involve two quantities. The first quantity, called the subgroup TPP ratio of $G$, is defined as
(1.3) $\rho_{0}(G):=\frac{\beta_{0}(G)}{|G|},$
where $\beta_{0}(G)$ is the subgroup TPP capacity of $G$ defined as
(1.4) $\beta_{0}(G):=\max\;\{|S||T||U|\mid(S,T,U)\textrm{ is a subgroup TPP triple of }G\}.$
The second quantity, called the TPP ratio of $G$, is defined as
(1.5) $\rho(G):=\frac{\beta(G)}{|G|},$
where $\beta(G)$ is the TPP capacity of $G$ defined as
(1.6) $\beta(G):=\max\;\{|S||T||U|\mid(S,T,U)\textrm{ is a TPP triple of }G\}.$
As subgroup TPP triples are special cases of TPP triples which are composed only of subgroups, clearly $\beta_{0}(G)\leq\beta(G)$ and $\rho_{0}(G)\leq\rho(G)$. Note that $G$ always realises the trivial subgroup TPP triple $(G,\{1\},\{1\})$, so that $\beta_{0}(G)\geq|G|$, or equivalently, $\rho_{0}(G)\geq 1$.
It is natural to seek best upper bounds for $\beta(G)$ and $\beta_{0}(G)$ (equivalently, for $\rho(G)$ and $\rho_{0}(G)$). It was shown by Cohn and Umans that if $G$ is a dihedral group then $\rho(G)\geq\frac{4}{3}$ [3]. P. Neumann derived a general upper bound for $\beta(G)$ [7, Corollary 3.2] that
(1.7) $\beta(G)\leq\left(\frac{1+\sqrt{1+8|G|}}{4}\right)^{3}<|G|^{3/2}$
using the following observation [7, Observation 3.1]:
Observation 1.2.
If $(S,T,U)$ is a TPP triple of a group $G$ then $|S|(|T|+|U|-1)\leq|G|$.
In 2012 Hedtke and Murthy conjectured [5, Table 1 and Conjectures 7.5-7.6], based on tables of TPP triple data in groups with cyclic subgroups of index $2$ of order up to $32$, obtained via exhaustive search algorithms in these groups implemented using the GAP computer algebra system, that
Conjecture 1.3.
If $G$ is a group with a cyclic subgroup of index $2$ then $\rho(G)\leq\frac{4}{3}$.
The author is aware of a proof (from a private communication from another researcher) of this conjecture for dihedral groups $D_{2n}$, i.e. that $\rho(D_{2n})\leq\frac{4}{3}$.
This note focuses on subgroup TPP ratio $\rho_{0}$ (equivalently, subgroup TPP capacity $\beta_{0}$), and proves a generalisation of the conjecture for $\rho_{0}$ for all groups with abelian normal subgroups of prime index. First, a few relevant properties and characterisations of TPP triples are stated, followed by proofs of two technical results needed to establish the main result. Following the proof of the main result some implications for $\rho$ for groups with cyclic normal subgroups of prime index are discussed, based on known data for $\rho$ in such groups.
2. Elementary Properties of TPP Triples
Let $G$ be a group.
Definition 2.1.
[7, p. 234] A TPP triple $(S,T,U)$ of $G$ is called basic if
(2.1) $S\cap T\cap U=\{1\}.$
For non-empty subsets $X,Y\subseteq G$ the quotient set $Q(X,Y)$ of $X$ and $Y$ is the subset $Q(X,Y)=XY^{-1}=\{xy^{-1}|x\in X,y\in Y\}\subseteq G$. If $X=Y$ then $Q(X,Y)=Q(X)=XX^{-1}$.
Theorem 2.2.
[5, Theorem 3.1] Three non-empty subsets $S,T,U\subseteq G$ satisfy the TPP if and only if
(2.2) $Q(S)\cap Q(T)Q(U)=Q(T)\cap Q(U)=\{1\},$
And if $S,T,U$ happen to be subgroups then the defining relation above simplifies to
(2.3) $S\cap TU=T\cap U=\{1\}.$
The following result states two invariance properties of TPP triples under permutations or certain kinds of set translations of its members.
Observation 2.3.
Let $(S,T,U)$ be any TPP triple of $G$, not necessarily basic in the sense defined above.
  1. (1)
    If $\pi\in S_{3}$, i.e. $\pi$ is a permutation of $\{S,T,U\}$, then $(S^{\pi},T^{\pi},U^{\pi})$ is a TPP triple of $G$ (permutation invariance) [3, Lemma 2.1].
  2. (2)
    If $a,b,c,d\in G$ are any elements then $(dSa,dTb,dUc)$ is a TPP triple of $G$ (translation invariance) [7, Observation 2.1].
The translation invariance property for TPP triples means that any non-basic TPP triple can be translated to a basic TPP triple of the same type and size as the original triple. To be precise, if $(S,T,U)$ is any non-basic TPP triple of $G$, that is, when $1\notin S\cap T\cap U$, then elements $s\in S$, $t\in T$, $u\in U$ can be chosen such that the right-translated triple $(Ss^{-1},Tt^{-1},Uu^{-1})$ is a basic TPP triple of $G$ with parameters $|Ss^{-1}|=|S|$, $|Tt^{-1}|=|T|$, $|Uu^{-1}|=|U|$. This means that generally only basic TPP triples need be considered. Of course, subgroup TPP triples are necessarily basic.
All further references to TPP triples will be to basic TPP triples, unless otherwise stated.
The following two elementary properties are useful to note, and will be used at several points.
Observation 2.4.
Given a TPP triple $(S,T,U)$ of a group $G$ and a subgroup $H\leq G$ there is a TPP triple $(S\cap H,T\cap H,U\cap H)$ of $H$.
Proposition 2.5.
[3, Lemma 3.1] Let $(S,T,U)$ be a TPP triple of $G$.
  1. (1)
[…]
  2. (2)
    If $G$ is abelian then the mapping $(s,t,u)\longmapsto s^{-1}tu$ on $S\times T\times U$ is injective into $G$, and $|S||T||U|\leq|G|$.
Part (2) of Proposition 2.5 means that only non-abelian groups can realise TPP triples of non-trivial size.
3. Technical Results
Some technical results needed for the main result are stated and proved. These are based on the idea of decomposing a TPP triple of a given group into smaller TPP triples obtained by independently decomposing the members of the triple with the left (or right) cosets of a suitable subgroup, using the set-closure property of TPP triples in relation to subgroups (see Observation 2.4).
Definition 3.1.
Let $G$ be a group, $H\leq G$ and $G/H=\{gH\mid g\in G\}$ the quotient group of $G$ by $H$ whose order is denoted by $|G:H|$ (also, $[G:H]$). For a non-empty subset $S\subseteq G$ let the $H$-support of $S$ be the set $\overline{S}:=\{gH\in G/H\mid S\cap gH\neq\emptyset\}\subseteq G/H$, i.e. the set of all cosets $gH\in G/H$ which intersect with $S$. If $(S,T,U)$ is a TPP triple of $G$ let its $H$-restriction be the triple $(S_{0},T_{0},U_{0})$, where $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, $U_{0}:=U\cap H$, and this is a TPP triple of $H$.
Observation 3.2.
Let $G$ be a group and $H\leq G$.
(1) If $S\leq G$ and $\overline{S}$ is its $H$-support, then $|\overline{S}|=|S:S\cap H|$.
(2) If $S\leq G$ and $H\trianglelefteq G$, i.e. $H$ is normal in $G$, then $\overline{S}\leq G/H$.
(3) If $H$ is abelian and $(S,T,U)$ is a subgroup TPP triple of $G$ then $|S||T||U|\leq\frac{\sigma\tau\upsilon}{n}|G|$, where $|G:H|=n$, and $\sigma=|S:S\cap H|$, $\tau=|T:T\cap H|$, $\upsilon=|U:U\cap H|$.
Proof.
Let $G$ and $H$ be given as above and $|G:H|=n$.
(1) Let $S\leq G$ and $\overline{S}$ be its $H$-support as defined above. Let $gH\in\overline{S}$ and $S_{g}:=S\cap gH$. By definition $S_{g}\neq\emptyset$. If $x\in S_{g}$ then $|S_{g}|=|S\cap gH|=|xS\cap xH|=|x(S\cap H)|=|S\cap H|$, where $x(S\cap H)$ is a coset of $S\cap H\leq S$. As the cosets $gH\in G/H$ are disjoint so are the sets $S_{g}$, which thus partition $S$. If $\sigma=|\overline{S}|$ then $|S|=\sigma|S\cap H|$, i.e. $\sigma=|S:S\cap H|$. The same result holds if left cosets are replaced by right cosets.
(2) Let $S\leq G$. If $H\trianglelefteq G$ then using the Second Isomorphism Theorem $SH\leq G$ and $\overline{S}=SH/H\leq G/H$, where $SH/H\cong S/(S\cap H)$.
(3) Let $(S,T,U)$ be a subgroup TPP triple of $G$. Then by Observation 2.4 $(S\cap H,T\cap H,U\cap H)$ is a subgroup TPP triple of $H$. If $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, $U_{0}:=U\cap H$ then by part (1) above $\sigma=|S:S_{0}|$, $\tau=|T:T_{0}|$, $\upsilon=|U:U_{0}|$, and, as by assumption $H$ is abelian, by Proposition 2.5 it follows that $|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{\sigma\tau\upsilon}\leq|H|$, which shows that $|S||T||U|\leq\sigma\tau\upsilon|H|=\frac{\sigma\tau\upsilon}{n}|G|$. ∎
Lemma 3.3.
Let $G$ be a group, $H\trianglelefteq G$ an abelian subgroup, $G/H=\{gH\mid g\in G\}$ the quotient group of $G$ by $H$, and $(S,T,U)$ a subgroup TPP triple of $G$. Furthermore, let the $H$-supports $\overline{S}$, $\overline{T}$, $\overline{U}$ be defined as in Observation 3.2, where these are subgroups of $G/H$.
(1) If $xH,yH$ are any non-trivial cosets of $H$ such that $S_{x},S_{y}\in\overline{S}$, $T_{x}\in\overline{T}$ and $U_{y}\in\overline{U}$, and letting $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, and $U_{0}:=U\cap H$, then $S_{x}^{-1}T_{x}U_{0}$ and $S_{y}^{-1}U_{y}T_{0}$ are non-trivial cosets of $S_{0}T_{0}U_{0}$ in $H$, that is, $S_{x}^{-1}T_{x}U_{0},S_{y}^{-1}U_{y}T_{0}\neq S_{0}T_{0}U_{0}$, and also distinct, that is, $S_{x}^{-1}T_{x}U_{0}\cap S_{y}^{-1}U_{y}T_{0}=\emptyset$.
(2) If $xH,yH$ are any non-trivial and distinct cosets of $H$ then $S_{x}^{-1}T_{x}U_{0}$ and $S_{y}^{-1}T_{y}U_{0}$ are distinct cosets of $S_{0}T_{0}U_{0}$ in $H$.
Proof.
Let $H$ be as given ($H$ is abelian and normal in $G$), and $H<G$ can be assumed, otherwise there is nothing to prove. Several facts used in the proof are noted for convenience:
  • The inverse of any coset $xH=Hx\in G/H$ is $(xH)^{-1}=x^{-1}H=Hx^{-1}$, and the inverse of any non-empty subset $B_{x}\subseteq xH$ is $B_{x}^{-1}\subseteq x^{-1}H=Hx^{-1}$ (by the normality of $H$ in $G$).
  • If $x_{1},x_{2}\in xH$ are any coset elements then $x_{1}x_{2}^{-1}\in H$ (if $x_{1}=xh_{1},x_{2}=xh_{2}$ for some $h_{1},h_{2}\in H$ then $x_{1}x_{2}^{-1}=xh_{1}h_{2}^{-1}x^{-1}=xx^{-1}h_{3}=h_{3}\in H$ for some $h_{3}\in H$ (by the normality of $H$ in $G$).
  • Every permutation $(S^{\pi},T^{\pi},U^{\pi})$ of the TPP triple $(S,T,U)$, for $\pi\in S_{3}$, is also a TPP triple (permutation invariance, Observation 2.3).
  • The subgroups $S_{0},T_{0},U_{0}$ as defined above, which satisfy the TPP, also have the property that $S_{0}\trianglelefteq S$, $T_{0}\trianglelefteq T$, $U_{0}\trianglelefteq U$ and $SH/H\cong S/S_{0}$, $TH/H\cong T/T_{0}$, $UH/H\cong U/U_{0}$ (Second Isomorphism Theorem). This means that translates of the form $S_{0}s_{x}=s_{x}S_{0},T_{0}t_{y}=t_{y}T_{0},U_{0}u_{z}=u_{z}U_{0}$, for $s_{x}\in S\cap xH$, $t_{y}\in T\cap yH$, $u_{z}\in U\cap zH$ and $xH,yH,zH\in G/H$, are cosets of $S_{0},T_{0},U_{0}$ respectively in $S,T,U$ respectively which intersect with the cosets $xH,yH,zH$ respectively.
  • As $H$ is abelian, the subgroups $S_{0}$, $T_{0}$, and $U_{0}$, are also normal in $H$ and thus form a set product $S_{0}T_{0}U_{0}$ which is a subgroup of $H$ of order $|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|\leq|H|$ (using the injective triple product map on $S_{0}\times T_{0}\times U_{0}$ into $H$ from Proposition 2.5).
(1) Let $xH\in\overline{S}\cap\overline{T}$ and $yH\in\overline{S}\cap\overline{U}$, and $S_{x},S_{y},T_{x},U_{y}$ be defined as above. Note that the sets $S_{x}^{-1}=(S\cap xH)^{-1}=S^{-1}\cap(xH)^{-1}=S\cap x^{-1}H$ and $S_{y}^{-1}=(S\cap yH)^{-1}=S\cap y^{-1}H$ and $S_{x}^{-1}T_{x},S_{y}^{-1}T_{y},S_{y}^{-1}U_{y}\subseteq H$ will also appear by calculation. As $S_{x},T_{x}\neq\emptyset$ elements $s_{x}\in S_{x}$, $t_{x}\in T_{x}$ can be chosen. Then, as $s_{x}^{-1}t_{x}\in H$ it follows that $s_{x}^{-1}t_{x}(S_{0}T_{0}U_{0})=S_{0}s_{x}^{-1}t_{x}T_{0}U_{0}=S_{x}^{-1}T_{x}U_{0}\subseteq H$, which shows that $S_{x}^{-1}T_{x}U_{0}$ is a coset of $S_{0}T_{0}U_{0}$ in $H$ of size $|S_{x}^{-1}T_{x}U_{0}|=|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|\leq|H|$. Additionally, if $xH\neq H$ (if $x\in G\backslash H$) then this is coset is non-trivial, i.e. $S_{x}^{-1}T_{x}U_{0}\neq S_{0}T_{0}U_{0}$. To see this, suppose there are elements $s_{x}\in S_{x}$, $t_{x}\in T_{x}$, $u_{0}\in U_{0}$ and $s_{0}\in S_{0}$, $t_{0}\in T_{0}$, $\widetilde{u}_{0}\in U_{0}$, such that $s_{x}^{-1}t_{x}u_{0}=s_{0}^{-1}t_{0}\widetilde{u}_{0}$. This can be rearranged as $s_{x}s_{0}^{-1}t_{0}\widetilde{u}_{0}u_{0}^{-1}t_{x}^{-1}=1$. As $t_{0}$ and $\widetilde{u}_{0}u_{0}^{-1}$ commute in $H$ this can be further rearranged as $s_{x}s_{0}^{-1}\widetilde{u}_{0}u_{0}^{-1}t_{0}t_{x}^{-1}=1$. However the TPP for $(S,U,T)$ implies that $s_{0}=s_{x}$ and $t_{0}=t_{x}$, a contradiction. By transposing $T$ and $U$, and $x$ and $y$, in the argument above, it can also be shown that $S_{y}^{-1}U_{y}T_{0}$ is a coset of $S_{0}T_{0}U_{0}$ in $H$, and, furthermore, a non-trivial coset if $yH\neq H$.
Suppose $xH,yH\neq H$ and $S_{x}^{-1}T_{x}U_{0}\cap S_{y}^{-1}U_{y}T_{0}\neq\emptyset$. Then there are elements $s_{x}\in S_{x}$, $s_{y}\in S_{y}$, $t_{x}\in T_{x}$, $u_{y}\in U_{y}$, $t_{0}\in T_{0}$, and $u_{0}\in U_{0}$ such that $s_{x}^{-1}t_{x}u_{0}=s_{y}^{-1}u_{y}t_{0}$, or $s_{x}s_{y}^{-1}u_{y}t_{0}u_{0}^{-1}t_{x}^{-1}=1$. Since $t_{0}$ and $u_{0}$ commute in $H$ this can be rewritten as $s_{x}s_{y}^{-1}u_{y}u_{0}^{-1}t_{0}t_{x}^{-1}=1$. However the TPP for $S,U,T$ implies that $u_{y}=u_{0}$ and $t_{x}=t_{0}$, a contradiction. So $S_{x}^{-1}T_{x}U_{0}\cap S_{y}^{-1}U_{y}T_{0}=\emptyset$ if $xH\in\overline{S}\cap\overline{T}$ and $yH\in\overline{S}\cap\overline{U}$ such that $xH,yH\neq H$, which proves part (1).
(2) Assume $U_{0}=U\cap H\trianglelefteq G$ and let $xH,yH\in\overline{S}\cap\overline{T}$ be such that $xh,yH\neq H$ and $xH\neq yH$. Suppose there are elements $s_{x}\in S_{x}$, $s_{y}\in S_{y}$, $t_{x}\in T_{x}$, $t_{y}\in T_{y}$, $u_{0},\widetilde{u}_{0}\in U_{0}$ such that $s_{x}^{-1}t_{x}u_{0}=s_{y}^{-1}t_{y}\widetilde{u}_{0}$, or, $st_{y}\bar{u}_{0}t_{x}^{-1}=1$, where $s=s_{x}s_{y}^{-1}$ and $\bar{u}_{0}=\widetilde{u}_{0}u_{0}^{-1}$. By the normality of $U_{0}$ in $G$ there is a $\hat{u}_{0}\in U_{0}$ such that $\bar{u}_{0}t_{x}^{-1}=t_{x}^{-1}\hat{u}_{0}$, which implies that $st_{y}\bar{u}_{0}t_{x}^{-1}=st_{y}t_{x}^{-1}\hat{u}_{0}=1$. However, the TPP for $S,T,U$ implies that $t_{x}=t_{y}$, a contradiction. Thus $S_{x}^{-1}T_{x}U_{0}$, $S_{y}^{-1}T_{y}U_{0}$ are distinct cosets of $S_{0}T_{0}U_{0}$ in $H$, which proves part (2). ∎
4. The Main Result
Theorem 4.1.
If $G$ is a group with an abelian normal subgroup of prime index $p$ then $\rho_{0}(G)\leq\frac{p^{2}}{2p-1}$.
Proof.
Let $G$ and $p$ be as given. Let $H\unlhd G$ be abelian of index $p$. Then $G/H$ is a cyclic group of order $p=|G:H|=\frac{|G|}{|H|}$ whose elements are $p$ (left) cosets $x_{0}H=H,x_{1}H,\ldots,x_{p-1}H$ of $H$ in $G$ with representatives $x_{0}=1,x_{1},\ldots,x_{p-1}\in G$ respectively. Let $(S,T,U)$ be a subgroup TPP triple of $G$ with $H$-supports $\overline{S},\overline{T},\overline{U}$ as defined in Observation 3.1. By part (2) of that Observation $\overline{S},\overline{T},\overline{U}\leq G/H$, and if $\sigma=|\overline{S}|$, $\tau=|\overline{T}|$, $\upsilon=|\overline{U}|$ then $\sigma,\tau,\upsilon\in\{1,p\}$. The sets $S_{x},T_{y},U_{z}$, as defined in Lemma 3.3 will also be used at certain points: these are defined as $S_{x}:=S\cap xH$, $T_{y}:=T\cap yH$, $U_{z}:=U\cap zH$ for any $xH,yH,zH\in G/H$. By definition, if $xH\in G/H$ then $xH\in\overline{S}\iff S_{x}\neq\emptyset$, and analagously for $T_{y}$ and $U_{z}$.
Let $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, $U_{0}:=U\cap H$. Then $(S_{0},T_{0},U_{0})$ is a subgroup TPP triple of $H$ (by Observation 2.4), and $S_{0}T_{0}U_{0}\mathrel{\unlhd}H$ (as $S_{0}T_{0}U_{0}\leq H$ and $H$ is abelian) and $|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{\sigma\tau\upsilon}\leq|H|$ and $|S||T||U|\leq\frac{\sigma\tau\upsilon}{p}|G|$ (by Proposition 2.5 and Observation 3.2).
Suppose $|S||T||U|=\sigma\tau\upsilon|S_{0}||T_{0}||U_{0}|>|G|=p|H|$. Then at least two of $\sigma,\tau,\upsilon$ must exceed $1$, that is, must be equal to $p$. Without loss of generality let $\sigma=\tau=p$, which means that $S$ and $T$ are supported on all $p$ cosets of $H$, or, equivalenty, $\overline{S}=\overline{T}=G/H$ and all sets $S_{x}$ and $T_{y}$ as defined above are non-empty. Suppose $\upsilon=1$, that is, $U=U_{0}$. Then $|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{p^{2}}>\frac{|H|}{p}$, i.e. $\frac{|H|}{|S_{0}||T_{0}||U_{0}|}=|H:S_{0}T_{0}U_{0}|<p$, which implies that $S_{0}T_{0}U_{0}=H$ and $|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|=|H|$, and, using Lemma 3.3, elements $s_{x}\in S_{x}\neq S_{0}$, $t_{x}\in T_{x}\neq T_{0}$ can be chosen so that the translate $s_{x}^{-1}t_{x}(S_{0}T_{0}U_{0})=S_{0}s_{x}^{-1}t_{x}T_{0}U_{0}=S_{x}^{-1}T_{x}U_{0}$ is a non-trivial coset of $S_{0}T_{0}U_{0}$ in $H$, a contradiction.
So, if $|S||T||U|>|G|=p|H|$ then $\sigma=\tau=\upsilon=p$ and $S,T,U\not\leq H$ and $|S_{0}T_{0}U_{0}|<|S_{0}||T_{0}||U_{0}|=\frac{|S||T||U|}{p^{3}}<|H|$. Also, as $H$ is maximal in $G$ and $U\not\leq H$ it follows that $U_{0}=U\cap H\trianglelefteq\langle U,H\rangle=G$ (and in the same way $T_{0}=T\cap H\trianglelefteq\langle T,H\rangle=G$). Now applying Lemma 3.3 shows that $\{S_{x}^{-1}T_{x}U_{0}\}_{xH\in G/H}$ and $\{S_{y}^{-1}U_{y}T_{0}\}_{yH\in G/H}$ are each collections of $p$ cosets of $S_{0}T_{0}U_{0}$ in $H$ having only the trivial coset $S_{0}T_{0}U_{0}$ in common. This means that $S_{0}T_{0}U_{0}$ has at least (2p1)(2p-1) cosets in $H$, i.e. $|H:S_{0}T_{0}U_{0}|=\frac{|H|}{|S_{0}||T_{0}||U_{0}|}\geq 2p-1$, which means that $|S||T||U|=p^{3}|S_{0}||T_{0}||U_{0}|\leq p^{3}\cdot\frac{|H|}{(2p-1)}=p^{3}\cdot\frac{1}{p(2p-1)}|G|$$\frac{p^{2}}{2p-1}|G|$.
A simple corollary is noted for some special cases,
Corollary 4.2.
If $G$ is a group with an abelian normal subgroup of prime index $p$ such that $(2p-1)\nmid|G|$, that is, (2p1)(2p-1) does not divide $|G|$, then $\rho_{0}(G)\leq\frac{1}{2}p$. If $G$ is a $p$-group with an abelian subgroup of index $p$ then $\rho_{0}(G)=1$.
Proof.
In the first case, if $H\leq G$ is abelian and normal of prime index $p$ such that $(2p-1)\nmid|G|$ then $(2p-1)\nmid|H|$, and if $(S,T,U)$ is a subgroup TPP triple of $G$ then the subgroup $S_{0}T_{0}U_{0}<H$, as defined above, cannot be as large as $\frac{|H|}{(2p-1)}$, i.e. $|S_{0}||T_{0}||U_{0}|<\frac{|H|}{(2p-1)}$, which means that $|S||T||U|=p^{3}|S_{0}||T_{0}||U_{0}|\leq p^{3}\cdot\frac{|H|}{2p}=\frac{1}{2}p|G|$. In the second case, if $G$ is a $p$-group then the product of the orders of any of its subgroups is a $p$-power, and if $G$ has an abelian subgroup of index $p$ (which is thus necessarily normal in $G$), and $(S,T,U)$ is a subgroup TPP triple of $G$, then $|S||T||U|$ must be a $p$-power such that $|S||T||U|<\frac{p^{2}}{2p-1}|G|<p|G|$, i.e. $|S||T||U|\leq|G|$. ∎
The contrapositive of this corollary explains the non-trivial $\rho_{0}$ values computed by Hedtke and Murthy for all the groups listed in [5, Tables 1-4] which satisfy the conditions of the corollary while achieving $\rho_{0}>1$ or $\rho_{0}>\frac{1}{2}p$, and this has been verified by GAP computations in these groups.
5. Analysis and Conclusion
In relation to the more general notion of TPP ratio $\rho$ the theorem means that in groups with abelian normal subgroups of prime index $p$ no three subgroups can realise the TPP with a triple size exceeding $\frac{p^{2}}{2p-1}|G|$. To achieve the latter one must look for triples of subsets at least one of which is not a subgroup.
In the case of $p=2$ and the dihedral groups $D_{2n}$ the theorem proves that $\rho_{0}(D_{2n})\leq\frac{4}{3}$ [5, Conjecture 7.5], and together with the data for TPP triples in groups of order $\leq 32$ obtained via exhaustive computational search [5, Table 1], strengthens the conjecture that $\rho\leq\frac{4}{3}$ for groups with cyclic (and normal) subgroups of index $2$ [5, Conjecture 7.6].
For $p>2$ the smallest known group containing a cyclic normal subgroup of index $p$ and a known $\rho$ is one of type $C_{7}\rtimes C_{3}$ (GAP ID [21, 1]), a non-abelian group of order $21=3\cdot 7$ containing a cyclic normal subgroup $C_{7}$ of index $p=3$, and which realises a largest TPP triple of type $(3,3,3)$ and size $3^{3}=27$, and satisfies the bound $\frac{27}{21}<\frac{3^{2}}{2\cdot 3-1}=\frac{9}{5}=1.8$ [5, Table 1]. The same data also shows that there are actually no groups of order $\leq 32$ containing cyclic normal subgroups of index $3$ and realising $\rho>\frac{9}{5}$. This seems to be true even for larger groups of order up to $100$, as indicated in the results of Xiang et. al. in 2018, who use a computational approach based on evolutionary search algorithms to look at TPP capacity $\beta$ (and other parameters related to group-theoretic matrix multiplication) in a selection of groups of order up to $100$ [8, Tables 1, 3-4]: some simple GAP computations show that of all groups listed in those tables which realise $\rho>\frac{9}{5}$ none contain cyclic normal subgroups of index $3$. Noting that a subgroup of a group which is of smallest prime index is always normal, this suggests the following conjecture.
Conjecture 5.1.
If $G$ is a group with a cyclic normal subgroup of prime index $p$, and $(S,T,U)$ is any TPP triple of $G$, then $|S||T||U|\leq\frac{p^{2}}{2p-1}|G|$, i.e. $\rho(G)\leq\frac{p^{2}}{2p-1}$. For this bound to be as tight as possible $p$ can be taken to be the smallest such prime, in which case the cyclic subgroup of index $p$ will necessarily be normal.
The conjecture is false for the more general case of abelian normal subgroups of prime index: for $p=2$, where the upper bound to be respected is $\rho\leq\frac{4}{3}$, the smallest known counterexamples are two $2$-groups of order $32$, one of type $(C_{4}\times C_{4})\rtimes C_{2}$ (GAP ID [32, 11]) and another of type $(C_{2}\times C_{2}\times C_{2}\times C_{2})\rtimes C_{2}$ (GAP ID [32, 27]), both containing a non-cyclic abelian subgroup of index $2$, but no cyclic abelian subgroups of index $2$, and realising $\rho=1.5>\frac{4}{3}$ (via TPP triples of type $(6,4,2)$).